How much heat is required to heat a 540g of iron from 300c to 65°c?
1
Expert's answer
2020-04-06T08:46:19-0400

The amount of heat needed to heat of iron sample

"Q=cm\Delta T""=0.450 \frac{\rm J}{\rm g\degree C}\times 540\:\rm g\times (65-30)\degree C=8505 \: J"